OpenStudy (anonymous):

Calculate the consumers' surplus at the indicated - unit price p for the demand equation. - p = 8 − 2q; p = 2

OpenStudy (anonymous):

is q the consumer surplus?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

so you are given the following equation. p = 8 − 2q you also know p = 2 so plug in 2 where you see p in the equation 2 = 8-2q subtract 8 from both sides. then divide by both sides to get an equation like q = something.